OpenStudy (anonymous):

I need help with solving a few rational equations using word problem form. I need to know how to work out the problem to

OpenStudy (kira_yamato):

Next one?

OpenStudy (anonymous):

I still havent figured out the first one.

OpenStudy (anonymous):

I don't mean to be wierd but do you by chance use yahoo or something cause this site keeps freezing and I am trying to finish my homework sooner rather than later.

OpenStudy (anonymous):

I just need some tutoring

OpenStudy (anonymous):

oh thank you

OpenStudy (katrinakaif):

9mph/0.25 + y = 6mph / y Cross multiply and get the time.

OpenStudy (anonymous):

ok I'm feeling lost again

OpenStudy (katrinakaif):

Well..You need to figure out what the time the daughter took.. Since the duaghter took 0.25 more hours than the dad the equation would be 0.25 + y y = How much hours the dad took Understood till here?

OpenStudy (anonymous):

somewhat

OpenStudy (anonymous):

So if I write it like this 9/.25+6/x, then cross multiply would that give me the distance of the trail?

OpenStudy (anonymous):

?

OpenStudy (katrinakaif):

In a proportion, we put the relative standards next to each other. So mph mph ---- = ----- time time Now plug it in 9mph 6mph ----- = ------ .25 + y y and then cross multiply

OpenStudy (anonymous):

ok

OpenStudy (katrinakaif):

Did you solve for y? (NOTE: y = time dad took

OpenStudy (anonymous):

I came out with .1666666666

OpenStudy (katrinakaif):

Hmm I came out with 0.5 When you cross multiply, you get 9y = 6(.25 + y) Distribute 6 9y = 1.5 + 6y 9y - 6y = 1.5 3y = 1.5 y = 0.5

OpenStudy (anonymous):

My instructor showed me to do cross mulitply then divide

OpenStudy (katrinakaif):

Hmm I am not familiar with that method..But my professor has taught me this way since I was in Middle School. =/

OpenStudy (anonymous):

for this problem the book says to use rt=d formula

OpenStudy (katrinakaif):

Yes we're getting to the point.

OpenStudy (anonymous):

oh ok

OpenStudy (katrinakaif):

Now since we have the dad's time which is 0.5, in order to find the daughters time, we must use the equation y + 0.25 0.5 + 0.25 = .75 which is the daughters time.

OpenStudy (katrinakaif):

O wait. Sorry. Its 0.5 - 0.25 Because daughter finished SOONER than her dad.

OpenStudy (katrinakaif):

So the daughter's time is 0.25 Now Use the equation RT = D R is the rate which in the daughter's case is 9mph T = 0.25 And Plug it in 0.25 ( 9) = D

OpenStudy (katrinakaif):

The distance is 2.25 units.

OpenStudy (katrinakaif):

Which in this case is miles ( as the rate used 'mph' or 'miles per hour' to express it)

OpenStudy (katrinakaif):

Do you understand?

OpenStudy (anonymous):

Kind of. I am looking back over what you wrote

OpenStudy (katrinakaif):

Review through it thoroughly and repeatedly =) Its fairly simple if you break it down step by step.